• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 11:39:11 on April28,2008

    Entry number 235470

    keyword=Target screen shots

    Here are some screen shots of the target's present status.

    We are at 2.5 µA beam current, 6 cm lH2 target.

    JP has just opened up the loop 2 (D2 loop) JT valve while closing
    down a bit the loop 1 (H2 loop) to cool down the deuterium loop,
    in order to get ready for some D2 data-taking which I understand
    is expected tomorrow.




    FIGURE 1

    FIGURE 2